OpenStudy (anonymous):

Meteorologist Wendy Stevens uses a theodolite (an angle-measuring device) on a 1-meter-tall tripod to find the height of a weather balloon. She views the balloon at a 44° angle of elevation. A radio signal from the balloon tells her that it is 1400 meters from her theodolite. a.How high is the balloon? b.How far is she from the point directly below the balloon? c.If Wendy’s theodolite were on the ground rather than on a tripod, would your answers change? Explain your reasoning.

OpenStudy (anonymous):

http://www.twiddla.com/527982

OpenStudy (dumbcow):

to find height height = 1400*sin(44)

OpenStudy (anonymous):

i got 972.5

OpenStudy (anonymous):

Answer is no. The thodolite is modelled as something very small, like a dot, so it doesn't matter on the ground or on the tripod.

OpenStudy (anonymous):

i get \[sin(44)=\frac{x}{1400}\]\[x=1400sin(44)=975.52\]but this is the height from one meter up so you have to add 1 to get 976.52

OpenStudy (anonymous):

for the second one if we call the distance d, then \[cos(44)=\frac{d}{1400}\] or \[d=1400cos(44)=1007\]

OpenStudy (anonymous):

this one we don't have to adjust.

OpenStudy (anonymous):

so for C the distance wouldn't change. i assume this question means if angle of elevation was still 44 degrees what would the height be? you would not have to add one to get the height and the distance would be the same

OpenStudy (anonymous):

it is yes or no

OpenStudy (anonymous):

it is not a yes or no question! distance in part b would be the same. for the height you would not have to add 1

OpenStudy (anonymous):

I answered your c question above. Same conclusion as satellite.